OpenStudy (anonymous):

Derivative and inverse function help!? Use the derivative to show why f(x)=definite integral 1/(4+e^t)dt, [0, ln x], with x>0 has an inverse function

OpenStudy (anonymous):

if f'(x) > 0 for all x in (0,inf) then f is strictly increasing, thus, it has an inverse. or if f'(x) < 0 for all x in (0,inf) then f is strictly decreasing, thus, it also has an inverse

OpenStudy (anonymous):

so after i find the derivative of this function, i look at the... range?

OpenStudy (anonymous):

what did you get for derivative

OpenStudy (anonymous):

i just realized i didn't put that the equation is 1/SQRT(4+e^t) but my derivative would be \[-\frac{ e^t }{ 2(e^t+4)^{3/2} }\]

OpenStudy (anonymous):

f'(x) = 1/[x sqrt(4 + x) ]

OpenStudy (anonymous):

again, f'(x) > 0 for all x in (0,inf) so f has an inverse

OpenStudy (anonymous):

f wasn't just 1/sqrt(4 + e^t) it was \[f(x) = \int\limits\limits_{0}^{\ln(x)} \frac{1}{\sqrt{4+e^t}} dt\]

OpenStudy (anonymous):

okay, wow, way to go wolfram derivatives... so because no matter what positive number is used for x, f'(x) is positive. would graphing help see the inverse?

OpenStudy (anonymous):

well, if you can find a way to graph f, then sure. But graphing is not always reliable

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=1%2F%28x%28sqrt%284%2Bx%29%29%29%2C+x+from+0+to+10 does this work as the graph of f'(x)?

OpenStudy (anonymous):

that one is f'(x), not f(x)

OpenStudy (anonymous):

but you can certainly use the graph of f'(x). As you can see, f'(x) is always positive on (0,inf)

OpenStudy (anonymous):

yeah, i graphed f ' (x). but do i need to figure out the inverse function? or is that not necessary?

OpenStudy (anonymous):

no, the question only asks for the proof that f(x) has an inverse.

OpenStudy (anonymous):

right, okay. so it has an inverse because the function is continuously increasing for all the values of x>0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

though I wouldn't say continuously increasing but rather strictly increasing.

OpenStudy (anonymous):

okay, strictly it is. Thank you so much!